# Feasibility study of True Muonium discovery with CERN-SPS H4 positron beam

### v2

@@ Abstract
True muonium ($\mu^{+}\mu^{-}$) is one of the heaviest and smallest electromagnetic bound states not containing hadrons, and has never been observed so far. In this work it is shown that the spin-1 TM state (ortho-TM) can be observed at a discovery level of significance in three months at the CERN SPS North-Area H4A beam line, using 43.7 GeV secondary positrons. In this way, by impinging the positrons on multiple thin low-Z targets, ortho-TM, which decays predominantly to $e^{+}e^{-}$, can be produced from $e^{+}e^{-}\to TM$ interaction on resonance ($\sqrt{s}\sim 2m_{\mu}$).
@@ I Introduction
  • at fixed-target experiments with positron beams of $\sim 43.7$ GeV (readily available at the CERN North-Area H4A beam line facility), to have $e^{+}e^{-}$ collisions at $\sqrt{s}\approx 2m_{\mu}$
@@
It is also possible to create TM from mesons decays:
  • $\eta\to TM\gamma$, with a branching ratio of $\sim 5\times 10^{-10}$. This production mode will be searched for at LHCb starting from the large $pp\to\eta X$ sample [6]
  • $K_{L}\to TM\gamma$, with a branching ratio of $\sim 7\times 10^{-14}$, searchable by future neutral-kaon based high-intensity experiments [7]
@@ IX Conclusion
In this paper, the authors explored the possibility of searching for TM in positron on target collisions at the CERN North-Area H4 beam line using a multi target approach. Each of the 10 target station includes 4 lithium targets followed by two very thin silicon detectors. According to the preliminary calculations and simulations described in this paper, a discovery can be obtained in few months of data taking.

### v3

@@ Abstract
True muonium ($\mu^{+}\mu^{-}$) is one of the heaviest and smallest electromagnetic bound states not containing hadrons, and has never been observed so far. In this work we show that the spin-1 TM state (ortho-TM) can be observed at a discovery level of significance in three months at the CERN SPS North-Area H4A beam line, using 43.7 GeV secondary positrons. In this way, by impinging the positrons on multiple thin low-Z targets, ortho-TM, which decays predominantly to $e^{+}e^{-}$, can be produced from $e^{+}e^{-}\to TM$ interaction on resonance ($\sqrt{s}\sim 2m_{\mu}$).
@@ I Introduction
  • at fixed-target experiments with positron beams of $\sim 43.7$ GeV (readily available at the CERN North-Area H4A beam line facility), to have $e^{+}e^{-}$ collisions at $\sqrt{s}\approx 2m_{\mu}$ [6]
@@
It is also possible to create TM from mesons decays [7, 8]:
  • $\eta\to TM\gamma$, with a branching ratio of $\sim 5\times 10^{-10}$ [9]. This production mode will be searched for at LHCb starting from the large $pp\to\eta X$ sample [10]
  • $K_{L}\to TM\gamma$, with a branching ratio of $\sim 7\times 10^{-13}$ [11], searchable by future neutral-kaon based high-intensity experiments
@@ IX Conclusion
In this paper, we explored the possibility of searching for TM in positron on target collisions at the CERN North-Area H4 beam line using a multi target approach. Each of the 10 target station includes 4 lithium targets followed by two very thin silicon detectors. According to the preliminary calculations and simulations described in this paper, a discovery can be obtained in few months of data taking.